OpenStudy (moongazer):

Calculate the force needed to hold a 2000-kg car on a ramp that makes an angle of 30 degrees with the horizontal, if the force is to be applied parallel to the ramp.

OpenStudy (anonymous):

resolving forces parallel to the slope, let force required = f and g = 9.8 f = 2000* g * sin 30 = 2000* 4.9 = 9,800 N i have assumed there is no friction between the car and the ramp.

OpenStudy (anonymous):

by the way moongazer i finally solved the question about the woman on the borad. the force on right hand support is 899 N so it would not be a safe siruation.

OpenStudy (moongazer):

yup your answer was correct :) thanks